Supose the perfectly competitve price is given as $46 and the total cost of the firm is given by TC=14X+2X^2, find



A) the prifit maximizing level of output for the firm



B) the profit of the firm



And show its graphical illustrations




1
Expert's answer
2022-01-04T11:12:24-0500

a) The profit maximizing level of output of the firm would be at the point where;

"P=MC"

"MC= dTC\/dx"

"MC=14+4x"

"P=46"

"46=14+4x"

"32=4x"

"x=8"

b) "Profit=TR-TC"

"TR=P.Q"

"TR=46*8 = 368"

"TC=14(8)+2(8*8) = 112+128 = 240"

"Profit =368 -240 =128"

Therefore Profit is $128.00